(19 Points) Caterpillars, Inc., a manufacturing company, acquired equipment on January 1, 2014 for $580,000. Estimated useful life of the equipment was seven years and the estimated residual value was $12,000. On January 1, 2017, after using the equipment for three years, the total estimated useful life has been revised to nine total years. Residual value remains unchanged. The company uses the straight-line method of depreciation. Calculate depreciation expense for 2017. (Round any intermediate calculations to two decimal places, and your final answer to the nearest dollar.) Select one O a. $54,095 O b. $63,111 О с. $64,444 O d. $55,238 (5 Points) A newly created design business, Rhodes Art, is finishing its first year of operations. During
Answer #1

Answer:

Correct answer is:

a. $54,095

Explanation:

Equipment acquired on 1st Jan 2014 for $580,000

Residual value = $12,000

Before Revision of useful life:

Useful life = 7 years

Depreciation per year under straight line method = (Cost - Residual value) / Useful life = ($580,000 - $12,000) / 7 = $81,142.86

Depreciation Accumulated depreciation Book Value $81,142.86 $498,857.14 162,285.72 $417,714.28 $243,428.58 $336,571.42 Year 2014 $81,142.86 2015 $81,142.86 2016 $81,142.86

As such book value of equipment at the end of 2016 or at beginning of 2017 is $336,571.42

On Revision of useful life:

Total estimated useful life has been revised to = 9 years

Remaining useful life at start of 2017 = 9 - 3 = 6 years

Residual value remains same.

Depreciation expense for 2017 = ($336,571.42 - $12,000) / 6 = $54,095

Hence:

Depreciation expense for 2017 = $54,095

As such option a is correct and other options b, c, d are incorrect.